OpenStudy (anonymous):

Write the equation of the line passing through the points (0, -4) and (-2, 2).

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

find the slope \[\frac{ y2 -y }{ x2 -x}\] then pick a point and use the equation \[y2-yo =m(x2-xo)\] your yo and xo are you co-ordinates. solve for y=mx+b

OpenStudy (anonymous):

usi\ng twjo